Determine if each ordered pair is a solution to the equation.
2x+y=4
a)(2,2)
b (8,-2)
I believe both are not an ordered pair to the equation. Am I right?

Answer by Alan3354(69443)

Determine if each ordered pair is a solution to the equation.
2x+y=4
a)(2,2)
b (8,-2)
I believe both are not an ordered pair to the equation. Am I right?
2x+y=4
a)(2,2)
Sub for x & y
2*2 + 2 =? 4
It doesn't.
-----------
b (8,-2)
2*8 - 2 =? 4
It doesn't.
---------
I believe both are not an ordered pair to the equation. Am I right?
I would say, "I believe neither is an ordered pair of the equation." But you are right.